MA114 Information for instructors Spring 2002 Course coordinator: Russell Brown Textbook: The textbook for this course will be Calculus, 3rd edition, by James Stewart, ISBN 0-534-21798-2 Syllabus: A course calendar with suggested problems is available. You may duplicate the version I have provided to Elizabeth. Each instructor should also provide an individual syllabus giving name, contact information and office hours and additional information about grading policy. The course calendar and a sample syllabus are available at http://www.ms.uky.edu/∼rbrown/courses/ma114cc.s.02 as tex documents. I expect that each instructor will prepare an additional page that details grading policies and gives their name, office hours etc. The University Ombudperson has made several requests for syllabi in all courses. 1. The syllabus should describe the curving policy for the course. This is apparently in response to instructors who used “negative curves”. 2. If attendance is to be used for grading, the syllabus should describe carefully what is meant by an excused absence. Please see the ombud’s memo for the full story. Math resource center: The mathematics resource center or Mathskeller opened in December 2001 in room 65 in the basement of the Classroom Building. All teaching assistants for MA114 are asked to schedule at least one of their office hours in this facility (and a total of three office hours). When the schedule is fixed, it will be circulated to all students in MA113 and 114. It is hoped that the extra work involved in holding office hours outside of the department will be balanced by having a quieter offices to work in the rest of the day. Pretest: Some (but not all) of our problems in calculus are due to the poor preparation of our students. Attached is a very simple pretest that you may choose to give in the first recitation. This will help to warn the students that they will be expected to know the basic facts from Calculus I in this course. Some students may be encouraged to drop, but at this point there is little they can do except retake MA113. Exams: There will be three exams and a final. These exams are scheduled in the evening at the times indicated in the common exam schedule. Though we have common exam times, we do not have common exams. Instructors in evening classes generally give their exams during a regularly scheduled class meeting. Rooms for exams will be assigned after classes begin. MA114 does not have a common final time. The final exam time is determined by the meeting time of the lecture, using the schedule in the Class Schedule Book or on the web, http://www.uky.edu/Registrar/final017.html MA194: In addition, to the 4 hours of credit for MA114, the department offers one additional hour of credit as MA194 on a pass/fail basis. Instructors may set their own grading policy for MA194. I suggest that students be passed in MA194 if they pass calculus and if they have no more than two unexcused absences. Thus, a student with three or more unexcused absences will fail. Below are a few common questions about MA194. Must a student take MA194? No, unless the student is in MathExcel. MathExcel students must take MA194. What section of MA194 should a student register in? Students should register in same section number for both MA194 and MA114. If a student drops or changes sections of MA114, they should also drop or change sections of MA194. Instructors and teaching assistants should check their MA194 rolls near the end of the semester. If there is a name that is unfamiliar, please try to determine if the student is registered in another section of MA114. If you have a student who thinks he or she is registered in MA194, but does not appear on your roll, try to find out if the student is registered in another section. Can MA194 be repeated? Yes, though there is little benefit to this. Can MA194 be taken without MA114? Yes, though it is not recommended. If the suggested grading is used, such students will fail MA194. Suggested grading for MA114: Students need 90% for an A, 80% for a B, 70% for a C and 60% for a D. Grades should be based on three hour exams, quizzes, homework, and longer assignments.

3 hour exams  300
Final exam  100
Homework, quizzes, 100
TOTAL  500

Homework: The list of problems on the syllabus is quite long. I plan to select a smaller number to assign to students. We will not have paper graders in Calculus II. Please do not burden your teaching assistants with excessive paper grading. Teaching assistants may be asked to grade short homework assignments or quizzes. Problems: Students should try to resolve problems with their instructor and/or teaching assistant. If this is not unsuccessful, see the departmental ombudperson, Ted Suffridge, or the university ombudperson. If you have complaints about the syllabus, please let me know. We will need a new course coordinator in the fall. Calculators: Most students are familiar with graphing calculators such as the TI-82. These calculators allow students to graph functions, solve equations, evaluate derivatives and definite integrals numerically. Elizabeth has TI-82 calculators that each instructor may check out. I suggest that students be allowed to use such calculators on exams. Test questions should be written so that it is clear whether a numerical answer from the calculator is acceptable or if students must carry out the computation by hand. In addition, some students will have machines that can carry out symbolic computations. I suggest that students not be allowed to use such machines on exams. Examples of such machines include the TI-89, TI-92, the HP48 and, of course, laptop computers.

Calculators will be useful in computing partial sums of series and drawing graphs of curves defined parametrically and in polar coordinates. Computer labs: The math department has a computer lab, Inslab, that instructors may use. See http://www.ms.uky.edu/∼inslab for more information. This lab have Maple and Matlab available. The university also maintains numerous computer labs where students may use Maple and other mathematical software. These labs contain classrooms that may be reserved.

Russell Brown

January 2, 2002